I need help on the following problem:
The difference when 7 less than a number is subtracted from twice the number is 12. What is the number?
I have no clue as to how to set up this equation. Please help!
thanks!

Answer by checkley71(8403)

2X-(X-7)=12
2X-X+7=12
X=12-7
X=5
PROOF
2*5-(5-7)=12
10-(-2)=12
10+2=12
12=12